OpenStudy (anonymous):

Is anyone taking Segment 1 of Algebra 2 on FLVS? Will Fan and Give Medal if anyone can help me !!!!!

Mehek (mehek14):

I'm taking Segment 1 of Algebra 1

OpenStudy (anonymous):

I really need help @Mehek14

Mehek (mehek14):

I didn't take Algebra 2

OpenStudy (anonymous):

I'm taking Algebra 2 and I really need help but I did Algebra 1. @Mehek14

OpenStudy (onepieceftw):

I'm in segment 2 of algebra two

OpenStudy (anonymous):

Did you take segment 1? @OnePieceFTW

OpenStudy (onepieceftw):

I did, what's your question? I'll help if I can.

OpenStudy (anonymous):

I don't understand one of the module's I need to do @OnePieceFTW

OpenStudy (onepieceftw):

Which one?

OpenStudy (anonymous):

01.05 and 03.06 are the two assignments I have to redo. I'm 61% done with the class those two are just bringing me down.

OpenStudy (onepieceftw):

What did you get wrong in 1.05?

OpenStudy (onepieceftw):

@yosoykiara ?

OpenStudy (anonymous):

My graphing was wrong. This is the last time I can submit it though. My friend was going to send me hers and show me how she did her graph but she deleted hers. @OnePieceFTW

OpenStudy (onepieceftw):

Launch Area:___(1, 2)___ Point A:___(0, 3)___ Point B:___(-3, 0)___ Point C:___(-1, -4)___

OpenStudy (onepieceftw):

Just put these values into geogebra, and there you have it.

OpenStudy (anonymous):

My teacher kept saying that there was more to it that was wrong which I'm not understanding

OpenStudy (onepieceftw):

That's weird... This is mine, but I don't remember what my grade on it was.

OpenStudy (anonymous):

I don't know my teacher kept saying it was wrong and my graph was wrong but I had the same graph you did.

OpenStudy (anonymous):

And also my 03.06 is messed up for some weird reason

OpenStudy (onepieceftw):

What did your teacher not like on that one?

OpenStudy (anonymous):

This is what she said: For #1 - What part of the quadratic formula can we use to determine the number and type of roots? What does this value tell us about the number and type of roots this function will have? #2 - Remember standard form is y=ax^2+bx+c, and the general form is y=a(x-h)^2+k. How did we learn in this lesson to convert from standard form to general form? (hint: what is the name of the lesson?) Remember to explain each step as you go. #3 - It looks like you have 2 sets of answers here. One of them is correct. Which one?

OpenStudy (onepieceftw):

For number one it should be . f(x) = x^2 + 6x – 16 Two real irrationals solutions g(x) = x^2 +6x + 1 Two real rationals solutions

OpenStudy (onepieceftw):

shoot, darn symbols

OpenStudy (onepieceftw):

F(x)=x^2+6x-16 G(x)=x^2+6x+1

OpenStudy (onepieceftw):

Okay so question 1 is just wanting you to identify the types of solutions. Question two is converting them to general vertex form. Three is actually solving them. (So sorry if I confused you, it' still coming back to me. :P)

OpenStudy (onepieceftw):

This is how I said in my submission to convert to vertex form: Write the equation substituting the f(x) for y. Add 16 to both sides. Divide 6 by two and square the result and add it to both sides of the equation to keep it balanced. Get the square root of the first and last term of the trinomial and write it like this: (x+3)^2 Subtract 25 from both sides. Result: y=(x+3)^2-25

OpenStudy (anonymous):

Sorry I'm a day late but thank you @OnePieceFTW